OpenStudy (bry856):

If f(x)=x^3−x, what is the average rate of change of f(x) over the interval [1, 5]? a)29.5 b)30 c)43 d)120

geerky42 (geerky42):

Generally, average rate of change of f(x) in interval [a, b] is \[\large \dfrac{f(b)-f(a)}{b-a}\]

geerky42 (geerky42):

So here you just have \(\dfrac{f(5)-f(1)}{5-1}=~\cdots~?\)

OpenStudy (bry856):

what, i dont know how to solve that im sorry

OpenStudy (anonymous):

@geerky42 did a good work

OpenStudy (rhystic):

you just plug in the 5 value into the equation you have subtract it by the equation with 1 plugged in and then put it over 5-1 which is 4.

OpenStudy (anonymous):

nothing left just put the numbers

OpenStudy (rhystic):

you're solving wrong. what did you get for 5^3-5 ?

OpenStudy (rhystic):

it should be ->( (5^3-5)-(1^3-1))/4